# Critical behavior of the Schwinger model via gauge-invariant VUMPS

### v1

@@ Abstract
We study the lattice Schwinger model by combining the variational uniform matrix product state (VUMPS) algorithm with a gauge-invariant matrix product ansatz that locally enforces the Gauss law constraint. Both the continuum and lattice versions of the Schwinger model with $\theta=\pi$ are known to exhibit first-order phase transitions for the values of the fermion mass above a critical value, where a second-order phase transition occurs. Our algorithm enables a precise determination of the critical endpoint in the continuum theory. We further analyze the scaling in the simultaneous critical and continuum limits and confirm that the data collapse aligns with the Ising universality class to remarkable precision.
@@ I Introduction
The phase structure and the critical phenomena in the Schwinger model have been the subject of much research. By semi-classical analysis and bosonization, Coleman [62] originally predicted phase transitions at $\theta=\pi$ for the fermion mass (divided by the dimensionful coupling $g$) above some critical value $(m/g)_{c}$. The existence of the phase transition was numerically confirmed by Hamer et al. [63], who diagonalized the lattice Hamiltonian for a small lattice and located the critical endpoint at $(m/g)_{c}=0.325\pm 0.02$. Schiller and Ranft [5] used the local Hamiltonian Monte Carlo method and obtained a comparable estimate. [^2] Concretely, [5] found $(m/g)_{c}=0.30(1)$ for $ga=1.0$ and $(m/g)_{c}$ $=0.31(1)$ for $ga=0.7$. Later, Byrnes et al. [12] utilized the DMRG method and found $(m/g)_{c}=0.3335(2)$. [^3] This result was confirmed by other works with different methods, including [20] (uMPS), [15, 16] (Grassmann tensor renormalization group), and [9, 10] (Monte Carlo method applied to the bosonized Schwinger model on the lattice). Our result obtained by gauge-invariant VUMPS is $(m/g)_{c}=0.333556(5)$.
In analyzing the ground state generated by the gauge-invariant VUMPS algorithm, the MPS transfer matrix is a useful quantity. In particular, it encodes information about the first excitation or equivalently the correlation length. We identify the critical endpoint of the lattice Schwinger model as the point in the parameter space where the correlation length diverges. Based on this, we determine the critical endpoint of the continuum theory by extrapolating to the zero lattice spacing.
@@
This paper is organized as follows. In Sec. II.1, we review the Kogut-Susskind formulation of the lattice Schwinger model. In Sec. II.2, we review the gauge-invariant uMPS ansatz for the ground state of the model. In Sec. III, we review the MPS transfer matrices. In Sec. IV, we determine the critical endpoint by two methods, namely, through the divergence of the correlation length in Sec. IV.1 and through the scaling analysis near the critical point in Sec. IV.2. We conclude the paper with discussion in Sec. V.
@@ II.1 Lattice-Regularized Schwinger Model
$\phi_{n}=\sigma_{-}(n)\prod_{l<n}(-i\sigma_{z}(l)).$ (6)
@@
$T|\ldots,s_{n},p_{n},\ldots\rangle=|\ldots,s_{n-1},p_{n-1},\ldots\rangle$ (9)
so that
$TL(n)T^{-1}=L(n+1)\,,\quad T\sigma_{i}(n)T^{-1}=\sigma_{i}(n+1)\,.$ (10)
@@
Since the mean expectation value of the electric field, $\langle\Psi|L(n)+L(n+1)+1|\Psi\rangle/2$ with $|\Psi\rangle$ being the ground state, is odd under the $CT$ transformation, it plays the role of a local order parameter.
@@ III Transfer Matrix and the correlation length
Here, $\xi$ represents the correlation length. $\eta$ takes $1/2$ with a small correction for one spatial dimension in a deep gapped phase, while it is determined by the argument of universality close to the critical point [74]. Comparing the exponential dependence of $n$ in the above expression with the one in eq. (21), the correlation length of the two-point function should be given by the smallest $\epsilon_{j}$ with $j\geq 1$ and $Z_{12}^{j}\neq 0$. Therefore the longest correlation length in the lattice unit is given by $1/\epsilon_{1}$.
We also define
@@
It was argued in [73] that $\delta(D)$ provides a useful parameterization of the deviation from the exact result due to the finiteness of the bond dimension $D$.
IV Locating the critical endpoint
We now present our simulation results and use them to determine the value of the critical endpoint $(m/g)_{c}$. The theta angle is fixed at $\theta=\pi$. With various input parameters $D$, $m/g$, and $ga$, we ran the gauge-invariant VUMPS simulations. Each run produced the matrices $(a_{n}^{s,q})_{\alpha_{q},\beta_{r}}$, from which we constructed the transfer matrix through (16) and (18). We then obtained the output parameters
@@ IV Locating the critical endpoint
The continuum limit corresponds to $ga\to 0$, while the thermodynamic limit corresponds to $\delta(D\to\infty)\to 0$.
@@ IV.1 Double extrapolations
In this subsection, we present our simulation results and use them to determine the critical mass $m_{c}$ in the continuum limit.
@@
Let us schematically describe the relations between the relevant quantities. For fixed values of $(ga\neq 0,m/g)$ and for large bond dimensions $D$, the two quantities $\epsilon_{1}$ and $\delta$ (both dependent on $D$) are approximately related as
@@
with $D$-independent parameters $\epsilon_{1,\infty}$ and $c_{1}$ (FIG. 1). The parameter $\epsilon_{1,\infty}=\epsilon_{1,\infty}(m/g)$ depends on $m/g$ as shown in FIG. 2. We recall from [63] that the lattice Schwinger model exhibits, even for a non-zero value of $ga$, a second-order phase transition as we vary $m/g$. Since $\epsilon_{1,\infty}$ is inversely proportional to the correlation length, it vanishes approximately linearly as $m/g$ approaches a ($ga$-dependent) critical value $(m/g)_{*}$:
@@
where $c_{-}$ and $c_{+}$ are $m/g$-independent positive parameters. The critical mass $m_{c}$ in the continuum limit can be obtained as a limit
@@
To numerically estimate $\epsilon_{1,\infty}=\epsilon_{1,\infty}(m/g,ga)$, $(m/g)_{*}$ $=m_{*}(ga)/g$, $(m/g)_{c}$, and their uncertainties, we proceed as follows. Let us take $G\equiv\{0.1,0.2,0.3,0.4\}$. For each $ga\in G$, we choose a set $S_{ga}$ of real numbers and we run VUMPS simulations for $m/g\in S_{ga}$ for the bond dimension $D$ large enough but not exceeding 500. For each $(ga,m/g)$, we fit the numerical data representing $(\delta,\epsilon_{1})$, obtained for high enough $D$, by the linear function (25) to obtain the value of $\epsilon_{1,\infty}$. In the standard linear fit where the data $(x_{i},y_{i})$ are fitted by a linear function $y=ax+b$, the uncertainty in $b$ can be estimated from the knowledge of the uncertainty (assumed uniform, i.e., independent of $i$) in the data $y_{i}$, and the value of the reduced chi-squared should be close to 1 for a good fit. In our case, we do not know the uncertainty in the value of $\epsilon_{1}$ for given $\delta$. We apply the common procedure [75] where we choose the uncertainty in $\epsilon_{1}$ such that the reduced chi-squared is 1, and use it to estimate the uncertainty in $\epsilon_{1,\infty}$. The uncertainties are shown as error bars in FIG. 2. We then fit the data set $\{(m/g,\epsilon_{1,\infty}(m/g))\,|\,m/g\in S_{ga}\}$ by a linear function in $m/g$ in each of the two regions $m/<(m/g)_{*}$ and $m/g>(m/g)_{*}$.555 Since we perform VUMPS for a finite number of $m/g$, when dividing the range of $m/g$ into the two regions, it is enough to know the value of $(m/g)_{*}$ only roughly and its $S$-dependence can be neglected. This fit gives the value of $m_{*}/g$ together with its uncertainty, for a given value of $ga$ and a choice of $S_{ga}$ (FIG. 2). Let us denote the tuple $(S_{ga=0.1},S_{0.2},S_{0.3},S_{0.4})$ by $S$. In the next step, we extrapolate to $ga=0$ by fitting $(m/g)_{*}$ with a polynomial in $ga$ and obtain the $S$-dependent estimate $m_{c,S}/g$ for the continuum critical mass and its uncertainty (FIG. 3).
@@
To quantify the sensitivity of $m_{c,S}/g$ to $S$ and the fitting method, we compute $m_{c,S}/g$ for many random choices of $S$ and two choices of the polynomial. We then randomly choose “good fits” of $(ga,(m/g)_{*})$, which we define to be those with the reduced chi-squared between 0.8 and 1.2. The resulting samples of $m_{c,S}/g$ are displayed in FIG. 4. The red-circle data points are obtained by fitting with a polynomial $C_{0}+C_{1}ga+C_{2}(ga)^{2}$ the data for $(ga,(m/g)_{*})$ obtained for $ga=0.1,0.2,0.3,0.4$ and various choices of $S$. For the estimate of $(m/g)_{c}$, we take the median value $M$ in this sample. As the estimate of the uncertainty in $(m/g)_{c}$, we take the sum of the median distance from $M$ and the median value of the uncertainty in $m_{c,S}$. Our final estimate for the continuum critical mass with uncertainty is
@@
Figure 4: Sorted values of $m_{c,S}/g$ computed for random choices of $S$ and two choices of fitting polynomials. In each data sequence, different data numbers correspond to different choices of $S$. The legend for for each sequence indicates the number of the values of $ga$ used to compute $m_{c,S}$ and the fitting polynomial.
@@ IV.2 Double collapse: A finite size scaling
In this subsection, we analyze the scaling behavior in the simultaneous critical and continuum limits and achieve the double data collapse method as discussed in [64, 65]. This provides further confirmation of our estimate (28).
@@
Here $\Delta_{t}^{\rm IR(UV)}$ and $\Delta_{\phi}^{\rm IR(UV)}$ are the critical exponents. For the entanglement entropy, we use the asymptotic formula based on the analysis of conformal field theory [77],
$S\sim\frac{c_{\rm UV}}{6}\log(\Lambda)+\frac{c_{\rm IR}}{6}\log(L),$ (34)
where $c_{\rm IR(UV)}$ is the central charge of conformal theories IR (UV) energy scales, and the lattice spacing is identified with $\Lambda^{-1}$. In the above expression, constant terms that do not depend on $\Lambda$ and $L$ are omitted.
Then scale invariant functions are given by
@@
$\tilde{S}[tL^{\Delta^{\rm IR}_{t}}\Lambda^{-\Delta^{\rm UV}_{t}}]=S-\frac{c_{\rm IR}}{6}\log(L)-\frac{c_{\rm UV}}{6}\log(\Lambda)$
Therefore, a randomly generated data set collapses into a single curve if all data points are sufficiently close to the critical endpoint.
Let us now identify $L=ga/\delta,~{}\Lambda=1/(ga),~{}\phi=\langle\Psi|L(n)+L(n+1)+1|\Psi\rangle/2,~{}\xi=1/\epsilon_{1}$ and $t=m/g-(m/g)_{*}$, where $(m/g)_{*}$ is the critical point of the lattice Schwinger model. $(m/g)_{*}$ is not universal and hence it is non-trivial function of $\delta$ and $ga$. We use the following ansatz of the critical point:
@@
We assume the Ising universality class, $\Delta^{\rm IR}_{m/g}=1,~{}\Delta_{\phi}=1/8,~{}c_{\rm IR}=1/2$, for the IR scale transformation, and $\Delta^{\rm UV}_{m/g}=0,~{}\Delta_{\phi}=0,~{}c_{\rm UV}=1$ for the UV scale transformation. The remaining parameters $(m/g)_{c},~{}b_{1}$ and $b_{2}$ should be optimized such that randomly generated data points are aligned into a single curve. Notice that nonzero $l_{1}$ yields universal shifts of $\tilde{\xi},~{}\tilde{\phi}$ and $\tilde{S}$.
In our analysis, the cost function is set to be the total length of all data points defined by
@@
where $\tilde{f}\equiv\tilde{\xi},~{}\tilde{\phi},~{}{\rm or}~{}\tilde{S}$, and $I$ represents the data point. The minimization problem is numerically solved by the gradient descent method. Due to the presence of many local minima near the global minimum, we solve the minimization problem multiple times with different initial guesses.
We randomly generate 1,780 data points near the critical endpoint within the ranges $0.3340.332<m/g<$ 0.334, $52050<D<$ 520, and $0.40.075<ga<$ 0.4. The best optimized values to achieve successful data collapse are determined for the correlation length, the order parameter, and the entanglement entropy as follows:
@@
All values are consistent with (28).
@@
FIG. 5 shows data points of scale invariant functions eq. (35) for the corresponding optimized parameter set. One can see from each panel that all data points collapse onto a single curve. The overall behavior of double collapse in FIG. 5 is the same as that of the $\lambda\phi^{4}$-theory [64, 65].

### v2

@@ Abstract
We study the lattice Schwinger model by combining the variational uniform matrix product state (VUMPS) algorithm with a gauge-invariant matrix product ansatz that locally enforces the Gauss law constraint. Both the continuum and lattice versions of the Schwinger model with $\theta=\pi$ are known to exhibit first-order phase transitions for the values of the fermion mass above a critical value, where a second-order phase transition occurs. Our algorithm enables a precise determination of the critical point in the continuum theory. We further analyze the scaling in the simultaneous critical and continuum limits and confirm that the data collapse aligns with the Ising universality class to remarkable precision.
@@ I Introduction
The phase structure and the critical phenomena in the Schwinger model have been the subject of much research. By semi-classical analysis and bosonization, Coleman [62] originally predicted phase transitions at $\theta=\pi$ for the fermion mass (divided by the dimensionful coupling $g$) above some critical value $(m/g)_{c}$. The existence of the phase transition was numerically confirmed by Hamer et al. [63], who diagonalized the lattice Hamiltonian for a small lattice and located the critical point at $(m/g)_{c}=0.325\pm 0.02$. Schiller and Ranft [5] used the local Hamiltonian Monte Carlo method and obtained a comparable estimate. [^2] Concretely, [5] found $(m/g)_{c}=0.30(1)$ for $ga=1.0$ and $(m/g)_{c}$ $=0.31(1)$ for $ga=0.7$. Later, Byrnes et al. [12] utilized the DMRG method and found $(m/g)_{c}=0.3335(2)$. [^3] This result was confirmed by other works with different methods, including [20] (uMPS), [15, 16] (Grassmann tensor renormalization group), and [9, 10] (Monte Carlo method applied to the bosonized Schwinger model on the lattice). Our result obtained by gauge-invariant VUMPS is $(m/g)_{c}=0.333556(5)$.
In analyzing the ground state generated by the gauge-invariant VUMPS algorithm, the MPS transfer matrix is a useful quantity. In particular, it encodes information about the first excitation or equivalently the correlation length. We identify the critical point of the lattice Schwinger model as the point in the parameter space where the correlation length diverges. Based on this, we determine the critical point of the continuum theory by extrapolating to the zero lattice spacing.
@@
This paper is organized as follows. In Sec. II.1, we review the Kogut-Susskind formulation of the lattice Schwinger model. In Sec. II.2, we review the gauge-invariant uMPS ansatz for the ground state of the model. We review the MPS transfer matrices in Sec. III. In Sec. IV, we determine the critical point by two methods, namely, through the divergence of the correlation length in Sec. IV.1 and through the scaling analysis near the critical point in Sec. IV.2. We conclude the paper with discussion in Sec. V.
@@ II.1 Lattice-Regularized Schwinger Model
$\phi(n)=\sigma_{-}(n)\prod_{l<n}(-i\sigma_{z}(l)).$ (6)
@@
$T|\ldots,s_{n},p_{n},\ldots\rangle=|\ldots,s_{n-1},p_{n-1},\ldots\rangle\,.$ (9)
@@
The mean expectation value of the electric field $\langle\Psi|(L(n)+L(n+1)+1)|\Psi\rangle/2$, where $|\Psi\rangle$ is a $T^{2}$-invariant ground state, is odd under the $CT$ transformation. Therefore, it serves as a local order parameter for $CT$-breaking.
@@ III Transfer Matrix and the correlation length
Here, $\xi$ is the correlation length in units of the lattice spacing $a$. The exponent $\eta$ takes the value $1/2$ with a small correction for one spatial dimension in a deep gapped phase [74, 75], while it is fixed by the scaling dimension close to the critical point [76]. Comparing the exponential dependence of $n$ in the above expression with the one in eq. (21), the correlation length of the two-point function should be given by the smallest $\epsilon_{j}$ with $j\geq 1$ and $Z_{12}^{j}\neq 0$. Therefore the longest correlation length in the lattice unit is given by $1/\epsilon_{1}$.
Reference [73] argued that the deviation from the exact result, arising from the finite bond dimension $D$, can be effectively parameterized by
@@
We will employ $\delta(D)$ in our analysis of the critical behavior of the lattice Schwinger model.
IV Locating the critical point
We now present our simulation results and use them to determine the value of the critical point $(m/g)_{c}$. The theta angle is fixed at $\theta=\pi$. With various input parameters $D$, $m/g$, and $ga$, we ran the gauge-invariant VUMPS simulations. Each run produced the matrices $(a_{n}^{s,q})_{\alpha_{q},\beta_{r}}$, from which we constructed the transfer matrix through (16) and (18). We then obtained the output parameters
@@ IV Locating the critical point
The continuum limit corresponds to $ga\to 0$, while the limit $\delta\to 0$ ($D\to\infty$) can be interpreted as removing the infrared cut-off [77, 64].
@@ IV.1 Double extrapolations
In this subsection, we determine the critical mass $m_{c}$ in the continuum limit via extrapolations, first in $\delta(D)$ and then in $ga$.
@@
Let us schematically describe the relations between the relevant quantities. For fixed values of $(ga\neq 0,m/g)$ and for large bond dimensions $D$, the two quantities $\epsilon_{1}$ and $\delta$ (both dependent on $D$) are approximately related as [^5] The assumption behind the ansatz (25) is that the finite bond dimension effect is similar to the finite size effect and discretizes the continuous spectrum. The quantity $\delta$ is a measure for how well the discrete spectrum approximates the exact continuous one [77, 64].
@@
with $D$-independent parameters $\epsilon_{1,\infty}$ and $c_{1}$ (FIG. 1). The parameter $\epsilon_{1,\infty}=\epsilon_{1,\infty}(m/g)$ depends on $m/g$ as shown in FIG. 2.666In FIG. 2, the data points at $m/g=0.3335$ and $0.3336$, which should be in the $CT$-unbroken phase according to the fitting lines for the other data points, are omitted. This is because for these values of $m/g$ the bond dimensions $D$ up to $500$ in our simulations are too small to generate a $CT$-invariant ground state, failing to show a straight line corresponding to the left half of the plot in FIG. 1. We recall from [63] that the lattice Schwinger model exhibits, even for a non-zero value of $ga$, a second-order phase transition as we vary $m/g$. Since $\epsilon_{1,\infty}$ is inversely proportional to the correlation length, it vanishes approximately linearly as $m/g$ approaches a ($ga$-dependent) critical value $(m/g)_{*}$:
@@
where $c_{-}$ and $c_{+}$ are $m/g$-independent positive parameters. The combination $CT$ of charge conjugation $C$ and translation $T$ is unbroken for $m<m_{*}$ and broken for $m>m_{*}$ [63], as can be confirmed from the expectation values of the local order parameter discussed at the end of Sec. II.1. The critical mass $m_{c}$ in the continuum limit can be obtained as a limit
@@
To numerically estimate $\epsilon_{1,\infty}=\epsilon_{1,\infty}(m/g,ga)$, $(m/g)_{*}$ $=m_{*}(ga)/g$, $(m/g)_{c}$, and their uncertainties, we proceed as follows. Let us take $G\equiv\{0.1,0.2,0.3,0.4\}$. For each $ga\in G$, we choose a set $V_{ga}$ of real numbers and we run VUMPS simulations for $m/g\in V_{ga}$ for the bond dimension $D$ large enough but not exceeding 500. For each $(ga,m/g)$, we fit the numerical data representing $(\delta,\epsilon_{1})$, obtained for high enough $D$, by the linear function (25) to obtain the value of $\epsilon_{1,\infty}$. In the standard linear fit where the data $(x_{i},y_{i})$ are fitted by a linear function $y=ax+b$, the uncertainty in $b$ can be estimated from the knowledge of the uncertainty (assumed uniform, i.e., independent of $i$) in the data $y_{i}$, and the value of the reduced chi-squared should be close to 1 for a good fit. In our case, we do not know the uncertainty in the value of $\epsilon_{1}$ for given $\delta$. We apply the common procedure [78] where we choose the uncertainty in $\epsilon_{1}$ such that the reduced chi-squared is 1, and use it to estimate the uncertainty in $\epsilon_{1,\infty}$. The uncertainties are shown as error bars in FIG. 2. We then fit the data set $\{(m/g,\epsilon_{1,\infty}(m/g))\,|\,m/g\in V_{ga}\}$ by a linear function in $m/g$ in each of the two regions $m/g<(m/g)_{*}$ and $m/g>(m/g)_{*}$. [^7] Since we perform VUMPS only for a finite set of the values of $m/g$, when dividing the range of $m/g$ into the two regions, it is enough to know the value of $(m/g)_{*}$ only roughly and its $V$-dependence can be neglected. This fit gives the value of $m_{*}/g$ together with its uncertainty, for a given value of $ga$ and a choice of $V_{ga}$ (FIG. 2). Let us denote the tuple $(V_{ga=0.1},V_{0.2},V_{0.3},V_{0.4})$ by $V$. In the next step, we extrapolate to $ga=0$ by fitting $(m/g)_{*}$ with a polynomial in $ga$ and obtain the $V$-dependent estimate $m_{c,V}/g$ for the continuum critical mass and its uncertainty (FIG. 3).
@@
To quantify the sensitivity of $m_{c,V}/g$ to $V$ and the fitting method, we compute $m_{c,V}/g$ for many random choices [^8] We fix a set $V^{(0)}_{ga}$ of real numbers, for example $V^{(0)}_{ga=0.1}=\{0.333$, $0.3331$, $0.3332$, $0.3333$, 0.3334}0.3334\}0.3334 }, and uniformly sample $V_{ga}$ from the set $\left\{V_{ga}\subseteq V_{ga}^{(0)}\,\middle|\,|V_{ga}|\geq 3\right\}$. of $V$ and two choices of the polynomial. We then choose “good fits” of $(ga,(m/g)_{*})$, which we define to be those with the reduced chi-squared between 0.8 and 1.2. The resulting samples of $m_{c,V}/g$ are displayed in FIG. 4, with the error bars indicating the uncertainties obtained by fitting [78]. The red-circle data points are obtained by fitting with a polynomial $C_{0}+C_{1}ga+C_{2}(ga)^{2}$ the data for $(ga,(m/g)_{*})$ obtained for $ga=0.1,0.2,0.3,0.4$ and various choices of $V$. For the estimate of $(m/g)_{c}$, we take the median value $M$ in this sample. As the estimate of the uncertainty in $(m/g)_{c}$, we take the sum of the median distance from $M$ and the median value of the uncertainty in $m_{c,V}$. Our final estimate for the continuum critical mass with uncertainty is
@@
The above fits show that the coefficient $C_{1}$ of $ga$ is rather small, motivating us to fit the data for $(ga,m_{*}/g)$ with $C_{0}+C_{2}(ga)^{2}$. This gives the blue-star ($ga=0.1,0.2,0.3,0.4$) and brown-cross ($ga=0.1,0.2,0.3$) data points shown in FIG. 4. The procedure in the previous paragraph gives estimates for $(m/g)_{c}$ as $0.333552(1)$ and $0.333554(2)$ for the data sets, respectively. These values are smaller than but consistent with (28).
Figure 4: Sorted values of $m_{c,V}/g$ computed for random choices of $V$ and two choices of fitting polynomials. In each data sequence, different data numbers correspond to different choices of $V$. The legend for for each sequence indicates the number of the values of $ga$ used to compute $m_{c,V}$ and the fitting polynomial.
@@ IV.2 Double collapse: A finite size scaling
In this subsection, we analyze the scaling behavior in the simultaneous critical and continuum limits of the lattice Schwinger model and achieve a double data collapse similar to the one observed for a different model in [64, 65]. This provides further confirmation of our estimate (28).
@@
Here $\Delta_{t}^{\rm IR(UV)}$ and $\Delta_{\phi}^{\rm IR(UV)}$ are the critical exponents. For the entanglement entropy, we assume the following scaling property,
$S\simeq\tilde{S}+\frac{c_{\rm UV}}{6}\log(\Lambda)+\frac{c_{\rm IR}}{6}\log(L),$ (34)
where $c_{\rm IR(UV)}$ is the central charge of conformal theories IR (UV) energy scales, and $\tilde{S}$ is a function invariant under the scale transformations (32). A contribution from a non-universal part is omitted. The scaling property with respect to $\Lambda$ is determined using the CFT formula presented in [79], where $\Lambda$ is identified as the inverse of the lattice spacing.
Then scale invariant functions are summarized as
@@
$\tilde{S}[tL^{\Delta^{\rm IR}_{t}}\Lambda^{-\Delta^{\rm UV}_{t}}]=S-\frac{c_{\rm IR}}{6}\log(L)-\frac{c_{\rm UV}}{6}\log(\Lambda).$
Therefore, a randomly generated data set collapses into a single curve if all the data points are sufficiently close to the critical point.
Let us now identify
$L$ $=\frac{ga}{\delta}\,,\quad\Lambda=\frac{1}{ga}\,,\quad t=\frac{m}{g}-\left(\frac{m}{g}\right)_{*}\,,$ (36)
$\phi$ $=\langle\Psi|(L(n)+L(n+1)+1)|\Psi\rangle/2,\quad\xi=\frac{1}{\epsilon_{1}}\,,$
where $(m/g)_{*}$ is the critical point of the lattice Schwinger model. $(m/g)_{*}$ is not universal and hence it is non-trivial function of $\delta$ and $ga$. We use the following ansatz of the critical point:
@@
We assume the Ising universality class, $\Delta^{\rm IR}_{t}=1,~{}\Delta_{\phi}=1/8,~{}c_{\rm IR}=1/2$, for the IR scale transformation, and $\Delta^{\rm UV}_{t}=0,~{}\Delta_{\phi}=0,~{}c_{\rm UV}=1$ for the UV scale transformation. The remaining parameters $(m/g)_{c},~{}b_{1}$ and $b_{2}$ should be optimized such that randomly generated data points are aligned into a single curve. Notice that nonzero $l_{1}$ yields universal shifts of $\tilde{\xi},~{}\tilde{\phi}$ and $\tilde{S}$.
In our analysis, we take as the cost function the total length of all data points defined by
@@
where $\tilde{f}\equiv\tilde{\xi},~{}\tilde{\phi},~{}{\rm or}~{}\tilde{S}$, and the index $I$ represents an ordered data point such that $\cdots<t_{I-1}L_{I-1}<t_{I}L_{I}<\cdots$. The minimization problem is numerically solved by the gradient descent method. Due to the presence of many local minima near the global minimum, we solve the minimization problem multiple times with different initial guesses.
We randomly and uniformly generate 1,780 data points near the critical point within the ranges $0.3340.332<m/g<$ 0.334, $52050<D<$ 520, and $0.40.075<ga<$ 0.4. The best-optimized values to achieve successful data collapse are determined for the correlation length, the order parameter, and the entanglement entropy as follows:
@@
All the values are consistent with (28).
@@
FIG. 5 shows data points of scale invariant functions eq. (35) for the corresponding optimized parameter set. One can see from each panel that all data points collapse onto a single curve. The overall behavior of double collapse in FIG. 5 is the same as that of the Euclidean $\lambda\phi^{4}$-theory [64, 65].